### v1

# Multilingual Performance of a Multimodal Artificial Intelligence System on Multisubject Physics Concept Inventories

Abstract
We investigate the multilingual and multimodal performance of a large language model-based artificial intelligence (AI) system, GPT-4o, on a diverse set of physics concept inventories spanning multiple languages and subject areas. The inventories taken from the PhysPort website cover the classical physics topics of mechanics, electromagnetism, optics, and thermodynamics as well as relativity, quantum mechanics, astronomy, mathematics, and laboratory skills. Unlike previous text-only studies, we uploaded the inventories as images mirroring what a student would see on paper, assessing the system’s multimodal functionality. The AI is prompted in English and autonomously chooses the language of its response – either remaining in the nominal language of the test, switching entirely to English, or mixing languages – revealing adaptive behavior dependent on linguistic complexity and data availability. Our results indicate some variation in performance across subject areas, with laboratory skills standing out as the area of poorest performance. Furthermore, the AI’s performance on questions that require visual interpretation of images is worse than on purely text-based questions. Questions that are difficult for the AI tend to be that way invariably of the inventory language. We also find large variations in performance across languages, with some appearing to benefit substantially from language switching, a phenomenon similar to code-switching of human speakers. Overall, comparing the obtained AI results to the existing literature, we find that the AI system outperforms average undergraduate students post-instruction in all subject areas but laboratory skills.
I Introduction
I.1 Generative artificial intelligence in physics education
The public availability of Large Language Models (LLMs), like those built on the architecture introduced by Vaswani et al. [1], has unlocked new possibilities across various domains, including education [2, 3]. Since the release of ChatGPT in the fall of 2022 [4], LLMs have surged in popularity, with scholars showcasing their remarkable capabilities. Beyond the wave of enthusiasm generated by human-like responses that have been shown to pass the Turing Test [5] with a majority of human test subjects [6], the GPT series demonstrated proficiency in academic fields such as physics [7]. Both the initial version and later iterations, particularly GPT-4 [8], have achieved impressive results, such as passing standardized exams, excelling in introductory courses and even coming close to passing entire degrees [9, 10, 11, 12, 13].
The technology is increasingly being embraced in physics education [14], offering promising applications for both teaching and learning physics. LLMs have proven valuable for teachers, helping in the creation of tailored materials and tasks, student assessments, and personalized feedback [15, 16, 17], as well as for their training [18]. For students, these models may represent ever-available, patient, and knowledgeable resources [19, 20]. However, these opportunities come with significant risks, particularly the potential for users to overly trust AI when assessing its scientific accuracy [21, 22].
Early publicly available generative AI systems have been limited to processing and generating text-based content only. Thus, earlier studies of AI’s performance on physics tasks, such as concept inventories, were restricted to text-based materials or textual descriptions of visual elements [9, 23]. More recently, multimodal systems, which can also input and output auditory and visual data, have broadened the scope of such studies (e.g., [24]).
In this study, we confront GPT-4o [25], a popular model from OpenAI, with screenshots from physics concept inventories, which reflect the test questions as seen by learners, including the accompanying diagrams, sketches, graphs, and illustrations. Earlier studies using GPT-4 and GPT-4o with the FCI [26], TUG-K [27] and BEMA [28] concept inventories demonstrated promising behavior, but also significant challenges, mainly related to the AI system’s limited ability to interpret visual information [29, 30, 31]. We considered the English version of each selected concept inventory and all their available translations.
I.2 Concept inventories in physics
Examinations in physics generally involve symbolic and numerical calculations, whereas concept inventories are typically different, focusing primarily on conceptual understanding. Scores on these different types of assessment do not necessarily correlate [32]. On the surface, this would work in AI’s favor, as AI systems have been notoriously “bad at math,” which hampered their performance on traditional physics exam questions [9]. However, early investigations of AI’s performance on conceptual tasks suggest that these, too, can present several challenges [33, 10].
Concept inventories have played and continue to play an important role in physics education research [34], and some of the most influential studies have been based on their outcomes, most notably with respect to active engagement [35]. Unlike traditional assessments that focus on individual learners, concept inventories are primarily designed to evaluate instructional methods, oftentimes with a focus on learning gains rather than the absolute scores. As we embark on assessing AI on the base of absolute scores, we deviate from this practice.
Arguably, while many concept inventory questions reflect students’ understanding of core ideas [36], they do not necessarily capture evidence of scientific practices or crosscutting concepts [37] (even though this distinction is debatable [38]). It is thus important to emphasize that our study will assess the equivalent of AI’s abilities and skills in physics, but not evaluate if it has the qualities of a physicist.
I.3 The language problem
Despite all the power and possibilities of using LLMs in education, there are also many limitations. OpenAI’s research on the GPT-4’s performance with different languages, for example, shows inconsistent results across languages [39]. LLMs learn patterns, make predictions, and perform tasks by leveraging large datasets, deep learning algorithms, and reinforcement learning from human feedback [40]. However, due to the disparities in the prevalence, quantity, and quality of information available across languages, there exists a disparity in the resources available for LLM training, which can have an impact on the quality of the outcomes. Nicholas and Bhatia highlight that although LLMs are designed to mitigate the issue of underrepresentation of certain languages in learning data, as of 2023, early LLMs were still predominantly trained on English materials. Although this may be changing as governments and companies attempt to strengthen AI capabilities in their countries (e.g., models like DeepSeek [41] or Qwen [42] developed by companies in China, or the Swiss AI Initiative [43]), it is likely that LLMs remain biased toward the needs of major economies with the financial resources to train AI systems. Thus, this can contribute to the transfer of cultural values and assumptions from English, or languages spoken in other major economies, to languages that lack infrastructure for training AI models. The authors stress the necessity to address this issue by companies, researchers, and governments across the globe [44]. Similar issues are also discussed in the analysis conducted by “Cohere for AI” company in their report titled “The AI Language Gap” [45]. Feng et al. also acknowledge the disparity in LLM performance across different languages, resulting in a gap of approximately 20% between high-resource and low-resource languages [46]. Particularly relevant for our study is a recent finding that a Chinese-trained model performed better on the FCI when prompted in Chinese rather than English [47].
Understanding and using physics-specific language is essential for physics literacy and learning. To address both the linguistic disparities of LLMs and the specialized nature of physics language, it is important to examine how LLMs perform across diverse languages in the context of physics education. While current research highlights the general challenges of multilingual use of LLMs, the intersection of language and subject-specific terminology, such as physics, is understudied and lacks a clear understanding of the current situation. This underscores the importance of evaluating LLMs not only in terms of language proficiency but also in their ability to accurately convey domain-specific knowledge.
Human speakers frequently overcome language disparities by engaging in code-switching [48, 49], switching languages mid-dialogue or even mid-sentence. Within physics education, much of the research on code-switching takes place in bilingual classrooms [50, 51], where students have a native language and a potentially different language of instruction; students tend to use their highest proficiency language when trying to understand what is happening conceptually [51].
I.4 Research questions
The overarching research questions are:
How does a state-of-the-art (2024/25) multimodal AI system perform across different physics concepts?
How does its performance compare to published post-instruction average results at the undergraduate level?
How does language influence the performance of the AI system?
How does the presence or necessity of images influence the performance of the AI system?
II Data set
As data set, we used concept tests published in PhysPort [52], which contains a comprehensive collection of research-based inventories in multiple translations [53]. We included tests that had at least a “bronze star” classification assigned on PhysPort, meaning that they had undergone significant research validation, and for which we could obtain an answer key. Figure 1 shows questions from the Force Concept Inventory in Persian and the Heat and Temperature Conceptual Evaluation in Chinese.
Table 1 shows the languages currently found in PhysPort, as well as the ISO 639 [54] shorthands we will be using throughout the paper; the rightmost column shows the number of inventories we were able to evaluate in each language. While the authors are familiar with some of the languages, the quality of the majority of the translations could not be evaluated.
Tables 3 through 6 show the concept tests included in our investigation, along with their available translations and literature references. For some inventories, more than one version was available. The column “%Post” lists some post-instruction inventory scores reported in the literature for undergraduate-level courses; these are collected best-effort and not necessarily representative (typically and much more systematically, inventory gains are reported [55]). We omitted results with very small sample sizes, as well as reported scores at the graduate-student or post-doctoral level. For subsequent comparisons, the scores found for each inventory were averaged.
The rightmost column “Cat.” shows the subject-area identifier under which we classified the test. Explanations for the abbreviations used in this column are provided in Table 2. These labels are derived from the PhysPort classification, however, we divided electricity and magnetism into EM-F (primarily dealing with fields and potentials), and EM-C (primarily focusing on DC and AC circuits). The multi-subject Next Gen Physical Science Diagnostic (NGPSD) was given no classification.
| Abbreviation | Description | Number |
| AR | Arabic | 4 |
| BN | Bengali | 1 |
| CA | Catalan | 1 |
| CS | Czech | 2 |
| DE | German | 5 |
| EL | Greek | 4 |
| EN | English | 53 |
| ES | Spanish | 16 |
| FA | Persian | 2 |
| FI | Finnish | 7 |
| FIL | Filipino | 1 |
| FR | French | 6 |
| HE | Hebrew | 2 |
| HI | Hindi | 1 |
| HR | Croatian | 1 |
| HU | Hungarian | 1 |
| ID | Indonesian | 5 |
| IS | Icelandic | 1 |
| IT | Italian | 2 |
| JA | Japanese | 9 |
| MS | Malay | 3 |
| NL | Dutch | 1 |
| NO | Norwegian | 1 |
| PA | Punjabi | 1 |
| PL | Polish | 1 |
| PT | Portuguese | 7 |
| RU | Russian | 1 |
| SK | Slovak | 1 |
| SR | Serbian | 1 |
| SV | Swedish | 8 |
| TA | Tamil | 1 |
| TH | Thai | 4 |
| TR | Turkish | 3 |
| UK | Ukrainian | 1 |
| ZH | Chinese | 8 |
| Abbreviation | Description |
| AST | Astronomy |
| EM-F | Electricity/Magnetism - Fields |
| EM-C | Electricity/Magnetism - Circuits |
| MATH | Mathematics |
| LAB | Laboratory Skills |
| MECH | Mechanics |
| OPT | Optics |
| QUANT | Quantum Physics |
| REAS | Reasoning |
| RELA | Relativity |
| THERM | Thermodynamics |
| Title | Full Title | Description | Refs. | Vers. | %Post | Languages | Cat. |
| ADT | Astronomy Diagnostic Test | Astronomy content knowledge (apparent motion of the sun, scale of the solar system, phases of the moon, linear distance scales, seasons, global warming, nature of light, gravity, stars, cosmology) | [56] | 2.0 |
41 [57]
54 [57] |
EN, ES, SV | AST |
| BEMA | Brief Electricity and Magnetism Assessment | Electricity/Magnetism content knowledge (circuits, electrostatics, magnetic fields and forces) | [28] | 1 |
42 [28]
43 [58] 61 [59] |
EN, ES, JA, PT, SV, ZH | EM-F |
| CCI | Calculus Concept Inventory | Mathematics content knowledge (functions, derivatives, limits, ratios, the continuum) | [60] | 5 |
50 [61]
52 [61] |
CS, EN | MATH |
| CDPA | Concise Data Processing Assessment | Lab skills (uncertainty in measurement, relationship between functions graphs and numbers) | [62] | 2 | 39 [62] | EN, ES | LAB |
| CSEM | Conceptual Survey of Electricity and Magnetism | Electricity / Magnetism content knowledge (electrostatics, magnetic fields and forces, Faraday’s law) | [63] | H |
61 [64]
66 [59] |
EN, ES, ID, MS, SV | EM-F |
| CTSR | Lawson Classroom Test of Scientific Reasoning | Scientific reasoning (proportional thinking, probabilistic thinking, correlational thinking, hypothetico-deductive reasoning) | [65] | 2 |
54 [66]
75 [66] |
EN, ES, SV | REAS |
| DIRECT | Determining and Interpreting Resistive Electric Circuit Concepts Test | Electricity / Magnetism content knowledge (DC circuits) | [67] | 1.2 |
44 [67]
63 [68] |
DE, EL, EN, ES, FI, SV, ZH | EM-C |
| DS | Density Survey | Mechanics content knowledge (density) | [69] | 1 | 57 [70] | DE, EN | MECH |
| ECA | Energy Concept Assessment | Mechanics content knowledge (energy principle, forms of energy, work and heat, absorption/emission spectrum, specifying appropriate systems) | [71] | 2 |
47 [71]
50 [71] |
EN, HR | MECH |
| ECCE | Electric Circuits Conceptual Evaluation | Electricity / Magnetism content knowledge (DC and AC circuits) | [72] | 1 |
37 [73]
42 [72] 64 [72] |
EN | EM-C |
| EMCA | Electricity and Magnetism Conceptual Assessment | Electricity / Magnetism content knowledge (electrostatics, electric fields and force, circuits, magnetism, induction) | [74] | 1 |
49 [74]
58 [74] |
EN, ID | EM-F |
| EMCS | Energy and Momentum Conceptual Survey | Mechanics content knowledge (energy, momentum) | [75] | 1 |
49 [75]
52 [75] 54 [76] 69 [76] |
EN, FI, ID, SV | MECH |
| FCI | Force Concept Inventory | Mechanics content knowledge (forces, kinematics) | [26] | v95 |
38 [77]
56 [78] 66 [79] |
AR, BN, CA, CS, DE, EL, EN, ES, FA, FI, FIL, FR, HE, HI, HR, HU, IS, IT, JA, MS, NL, NO, PA, PL, PT, RU, SV, SK, TA, TH, TR, ZH | MECH |
| FMCE | Force and Motion Conceptual Evaluation | Mechanics content knowledge (kinematics, forces, energy, graphing) | [80] | v99 | 55 [81] | EN, ES, ID, JA | MECH |
| FORT | Montana State University Formal Reasoning Test | Scientific reasoning (hypothesis testing, correlational reasoning, probability, control of variables, proportional reasoning) | [82] | 1 | 55 [82] | EN | REAS |
| FTGOT | Four-tier Geometrical Optics Test | Waves / Optics content knowledge (plane mirrors, spherical mirrors, lenses); used as two-tier test | [83] | 1 | 18 [83] | EN, TR | OPT |
| FVA | Force, Velocity, and Acceleration Test | Mechanics content knowledge (forces, velocity, acceleration) | [84] | 3.2.3a |
38 [84]
70 [84] |
EN | MECH |
| Title | Full Title | Description | Refs. | Vers. | %Post | Languages | Cat. |
| GECI | Greenhouse Effect Concept Inventory | Astronomy content knowledge (types of greenhouse gases, types of electromagnetic energy, energy equilibrium balance, greenhouse effect mechanisms, global warming vs. greenhouse effect) | [85] | vC | 55 [85] | EN, JA | AST |
| HTCE | Heat and Temperature Conceptual Evaluation | Thermal / Statistical content knowledge (temperature, phase change, heat transfer, thermal properties of materials) | [86] | 1 | 78 [86] | EN, ZH | THERM |
| IBCDC | Inventory of Basic Conceptions – DC Circuits | Electricity / Magnetism content knowledge (DC circuits) | [87] | F06 |
EN: 40
FR: 35 [87] |
EN, FR | EM-C |
| IBCM | Inventory of Basic Conceptions – Mechanics | Mechanics content knowledge (forces, kinematics) | [87] | F06 |
EN: 26
FR: 31 [87] |
EN, FR | MECH |
| LPCA | Light Phenomena Conceptual Assessment | Waves / Optics content knowledge (reflection, refraction, Snells law, wavelength and frequency, light scattering, electromagnetic spectrum, the human eye) | [88] | 1 | 41 [88] | EN | OPT |
| LPCI | Lunar Phases Concept Inventory | Astronomy content knowledge (phases of the moon) | [89] | 3 |
42 [89]
55 [89] |
EN, ES | AST |
| LSCI | Light and Spectroscopy Concept Inventory | Astronomy content knowledge (light, waves, spectroscopy) | [90] | 1 |
47 [91]
51 [91] 52 [91] |
EN | AST |
| MBT | Mechanics Baseline Test | Mechanics content knowledge (kinematics, forces, momentum, energy) | [92] | 97 |
35 [93]
48 [94] 66 [95] 73 [95] |
DE, EL, EN, ES, FA, FI, FR, IT, JA, MS, PT, TR | MECH |
| MCS | Magnetism Conceptual Survey | Electricity / Magnetism content knowledge (magnetic fields and forces, Faraday’s law) | [96] | 1 |
41 [96]
44 [96] |
EN | EM-F |
| MUQ | Measurement Uncertainty Quiz | Lab skills (calculating error from measurements, accuracy and precision, sources of error) | [97] | 1 | — | EN | LAB |
| MWCS | Mechanical Wave Conceptual Survey | Waves / Optics content knowledge (mechanical waves, wave propagation, wave superposition, wave reflection, standing waves) | [98] | 1 | 47 [99] | EN, ES, TH | OPT |
| 2 | EN, ES | ||||||
| NGCI | Newtonian Gravity Concept Inventory | Astronomy content knowledge (directionality of gravity, force law, thresholds related to gravity, independence of forces) | [100] | 3 |
50 [100]
55 [100] |
AR, EN | AST |
| NGPSD | Next Gen Physical Science Diagnostic | Mechanics content knowledge (magnetism, static electricity, energy, forces, waves and sound, light) | [101] | 2 | — | EN | — |
| PIQL | Physics Inventory of Quantitative Literacy | Mathematics scientific reasoning (proportional reasoning, reasoning with signed quantities, co-variational reasoning) | [102] | 2.2 | 55 [102] | EN | REAS |
| Title | Full Title | Description | Refs. | Vers. | %Post | Languages | Cat. |
| QMCA | Quantum Mechanics Concept Assessment | Modern / Quantum Content knowledge (wave functions, measurement, time dependence, probability, infinite square well, 1D tunneling, energy levels, spins) | [103] | 5.5.7 | 54 [103] | EN, PT | QUANT |
| 6.6.2 | EN | ||||||
| QMCS | Quantum Mechanics Conceptual Survey | Modern / Quantum content knowledge (wave functions, probability, wave-particle duality, uncertainty principle, infinite square well, one-dimensional tunneling, energy levels) | [104] | 2.0 |
51 [104]
69 [104] |
EN, FI, JA | QUANT |
| QMFPS | Quantum Mechanics Formalism and Postulates Survey | Modern / Quantum content knowledge (quantum mechanics formalism, quantum mechanics postulates) | [105] | 29 |
32 [106]
37 [106] |
EN, ES | QUANT |
| QMS | Quantum Mechanics Survey | Modern / Quantum content knowledge (wave functions, measurement, expectation values, Hamiltonian, time dependence, probability, infinite square well, finite square well, harmonic oscillator, 1D tunneling) | [107] | 18 | 38 [107] | EN | QUANT |
| QMVI | Quantum Mechanics Visualization Instrument | Modern / Quantum content knowledge (wave functions, probability, infinite square well, 1D tunneling, time dependence, momentum space, 2D potentials, visualization of the relationship between potentials and wave functions) | [108] | 0.4 |
28 [108]
29 [108] 45 [108] 58 [108] |
EN | QUANT |
| QPCS | Quantum Physics Conceptual Survey | Modern / Quantum content knowledge (de Broglie wavelength, double slit interference, uncertainty principle, photoelectric effect, wave particle duality) | [109] | 1 |
59 [109]
75 [109] |
EN, TH | QUANT |
| RAPT | Rate and Potential Test | Electricity / Magnetism Content knowledge (electric potential, rate of change) | [110] | 1A |
61 [110]
73 [110] |
EN | EM-F |
| RCI | Relativity Concept Inventory | Modern / Quantum content knowledge | [111] | 1 | 71 [111] | EN | RELA |
| RFCI | Representational Variant of the Force Concept Inventory | Mechanics content knowledge (kinematics, forces, graphing, multiple representations) | [112] | 2007 |
60 [112]
61 [112] |
EN, FI | MECH |
| RKI | Rotational Kinematics Inventory | Mechanics content knowledge (Part 1: rotational kinematics of a particle, Part 2: rotational kinematics of a particle in rectilinear motion, Part 3: rotational kinematics of a rigid body about a fixed axis) | [113] | 1 |
38 [114]
55 [114] |
EN, FR | MECH |
| RRMCS | Rotational and Rolling Motion Conceptual Survey | Mechanics content knowledge (rotational kinetic energy, torque, rotational kinematics, moment of inertia) | [115] | 1 | 75 [115] | EN | MECH |
| Title | Full Title | Description | Refs. | Vers. | %Perc | Languages | Cat. |
| SGCE | Symmetry and Gauss’s Law Conceptual Evaluation | Electricity / Magnetism content knowledge (symmetry, electric field, electric flux) | [116] | 1 | 49 [116] | EN | EM-F |
| SPCI | Star Properties Concept Inventory | Astronomy content knowledge (stellar properties, nuclear fusion, star formation) | [117] | 4 |
30 [117]
51 [117] |
EN, ES, JA | AST |
| STPFaSL | Survey of Thermodynamic Processes and First and Second Laws | Thermal / Statistical content knowledge (first law of thermodynamics, second law of thermodynamics, PV diagrams, reversible processes, irreversible processes) | [118] | short | 37 [118] | EN, ID, PT, ZH | THERM |
| long | EN | ||||||
| TCE | Thermal Concept Evaluation | Thermal / Statistical content knowledge (temperature, heat transfer, phase change, thermal properties of materials) | [119] | 1 | 78 [119] | EN, JA, PT, ZH | THERM |
| TCS | Thermodynamic Concept Survey | Thermal / Statistical content knowledge (temperature, heat transfer, ideal gas law, first law of thermodynamics, phase change, thermal properties of materials) | [120] | 2 |
43 [120]
46 [120] 58 [120] |
EN, TH, ZH | THERM |
| TOAST | Test of Astronomy Standards | Astronomy content knowledge (gravity, electromagnetic radiation, fusion and formation of heavy elements, evolution of the universe, star and stellar evolution, evolution and structure of the solar system, seasons, scale, yearly patterns, daily patterns, moon phases) | [121] | vf | 44 [121] | EN, JA | AST |
| TUG-K | Test of Understanding Graphs in Kinematics | Mechanics content knowledge (kinematics, graphing) | [27] | 2.6 | 59 [122] | AR, DE, FI, FR, HE | MECH |
| 3.0-4.0 | EL, EN, ES, PT, SV, UK, ZH | ||||||
| TUV | Test of Understanding of Vectors | Mathematics content knowledge (magnitude, direction, components, unit vector, addition, subtraction, multiplication, dot and cross product) | [123] | 1 | 68 [123] | AR, EN, ES | MATH |
III Methodology
III.1 Data preparation
The questions from the concept tests were captured using screenshots like the ones shown in Fig. 1. If a question had multiple parts referring to the same scenario or each other, these were combined in one image; at times, this required manual photo-editing to close page breaks. Questions that assessed confidence levels (“How sure are you of this answer?”) or those with free-response answers were skipped. The tests in Tables 3–6 resulted in 3,662 images.
The solution keys were transcribed from PhysPort, where we transformed the various option values such as B), 2., b., ), and non-Arabic, non-Latin, non-Greek characters into lower-case Latin characters. For tests that gave partial credit for certain incorrect answers if previous answers were incorrect (“follow-up mistakes”), we simplified the evaluation by only giving credit for correct answers. Also for tests that assigned different point values for different questions, we only considered correctness. In addition, for each question, we coded if it was text-only, or if it included an image, graph, or scenario illustration.
III.2 AI processing
We have been using GPT-4o [25] Version 2024-08-06 via Microsoft Azure AI Services [124] at ETH Zurich. The university’s contract includes provisions that any data submitted will not be used for training purposes; this provision is crucial to avoid compromising the confidentiality and validity of the concept tests.
The screenshots were submitted via the deployment’s API in base64-encoding as shown in Fig. 2. As role, we used the text shown in Fig. 3, and as prompt the text shown in Fig. 4. To facilitate further processing and evaluation of the answers, we demanded the structured JSON output schema shown in Fig. 5. We included the “explanation” field in the output structure to capture the AI system’s reasoning steps, also referred to as Chain-of-Thought (CoT) [10]. This allowed us to glean some key insights into the system’s language behavior.
Each screenshot was independently evaluated three times at a temperature of 0.7 (which is the default in the chat clients), and the resulting three problem structures were combined into a solution array. This resulted in output like the one shown in Fig. 6. Altogether, we obtained 14022 solutions for 4,674 questions (1,498 for English and 3,176 for non-English language inventories).
You are a physics and mathematics expert. You are given images of multiple-choice test questions, which you will answer correctly in JSON format. You work very carefully, and you strongly favor correct answers over rapid responses. You do mathematical calculations and derivations step-by-step. If there are graphics, you consult them more than once if needed in order to get information for your reasoning.
You are tasked with correctly solving the problem or problems from the image. The image should contain problem(s) with the problem identifiers: [Problems] For this multiple-choice problem or these multiple-choice problems, carefully consider the situation and document your reasoning, and only then pick the answer choice that aligns with your reasoning.
As Fig. 6 illustrates, the role and prompt we used can result in mixed-language outputs. While this issue could probably have been avoided by translating the role and prompts into the language of the tests, doing so for languages using non-ASCII characters would likely have made the mapping between the prompt and the fields in the JSON structure unreliable. Colloquially speaking, “the AI was allowed to think in the language of its choice.”
III.3 Analysis methods
The answer choices found in the JSON structures were normed to the same lower-case Latin characters as the solution keys. In cases where the AI did not provide a valid answer (e.g., where it claimed that there was no correct answer specified), the answer was counted as incorrect. Some concept inventories had images or scenarios labeled with Roman characters and answer choices such as “A) I, B) II, C) III, D) IV, E) none of the above;” in this case, if the AI picked “IV,” this was manually converted to “d.” Another source of possible error were numbered scenarios within numbered multipart problems; in this case, the AI at times ignored the problem numbers provided in the prompt and instead used the scenario numbers; this was fairly obvious during the evaluation, but had to be fixed manually.
We counted each answer in an inventory — three answers per question — as either correct or incorrect, and considered the percentage of correct answers as performance measures. For each answer, we coded the language of both the test description and the answer explanation, primarily using langdetect [125] and manual determination in some cases. Responses were categorized as fully in the language of the test, fully in English, or mixed. The most common mixed scenario occurred when the problem description was still in the language of the test, but the explanation was in English. However, there were also rare cases where the language switched mid-stream for either of those.
III.4 Use of AI
While obviously being the subject of this study, AI (GPT-o1 [126]) has also been used for the following aspects of the study:
Initial drafts of analysis programs in R and Python.
Exploratory data analysis.
LaTeX formatting of manuscript components.
Improving the grammar and readability of manuscript passages.
IV Results
IV.1 Scores by inventory and language
Table 7 shows the scores on each of the inventories in each available nominal language (as “nominal,” we denote the language in which the inventory was presented). One immediate observation is the uneven coverage of assessments across languages. Many inventories are only available in a handful of languages, making broad comparisons challenging. However, when data is present, certain patterns do emerge. Some assessments, such as FCI, have results for numerous languages, allowing a more comprehensive look at language-to-language differences.
In the FCI, performance ranges from as low as 20% in Punjabi and 22% in Tamil to as high as 74% in Portuguese and Polish, suggesting that the AI’s proficiency can vary significantly depending on the language used. Since the FCI uses five answer options, a 20% performance is equivalent to randomly picking an answer. An in-depth analysis of the FCI answers for Punjabi revealed that GPT-4o started hallucinating, generating output about uniform magnetic fields and charged particles. In Tamil, the explanations and conclusions were sometimes correct, but then GPT-4o picked the wrong answer letter. Other assessments, like TUV, ADT, and FMCE, show relatively consistent scores in the 60–80% range wherever data is available.
A second key point is that some inventories appear to have stable performance across multiple languages, while others do not. For example, QMCS shows high and relatively stable performance (78–86%) in languages for which data is available, indicating that the AI handles that conceptual domain and its translations fairly well. In contrast, MBT scores fluctuate more dramatically (e.g., 27% in Persian, 53% in Finnish, and 44% in Italian and Hungarian), indicating that performance may depend not just on the conceptual domain but also on the subtleties of the language and phrasing used.
One such outlier is the TUG-K2.6, where performance differs widely across the few languages tested, from 48% in Arabic to 67% in Finnish, with little intermediate data. Similarly, TUG-K3.0-4.0 shows a scattered pattern: 44% in Greek, 55% in English, 60% in Spanish, and dips to 45% or lower in other languages, suggesting highly variable understanding.
| AR | BN | CA | CS | DE | EL | EN | ES | FA | FI | FIL | FR | HE | HI | HR | HU | ID | IS | IT | JA | MS | NL | NO | PA | PL | PT | RU | SK | SR | SV | TA | TH | TR | UK | ZH | |
| ADT | - | - | - | - | - | - | 78 | 81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 83 | - | - | - | - | - |
| BEMA | - | - | - | - | - | - | 66 | 49 | - | - | - | - | - | - | - | - | - | - | - | 53 | - | - | - | - | - | 57 | - | - | - | 51 | - | - | - | - | 44 |
| CCI | - | - | - | 79 | - | - | 83 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CDPA | - | - | - | - | - | - | 33 | 40 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CSEM | - | - | - | - | - | - | 54 | 53 | - | - | - | - | - | - | - | - | 38 | - | - | - | 53 | - | - | - | - | - | - | - | - | 45 | - | - | - | - | - |
| CTSR | - | - | - | - | - | - | 86 | 74 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 62 | 76 | - | - | - | - | - |
| DIRECT | - | - | - | - | 45 | 53 | 49 | 33 | - | 52 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 49 | - | - | - | - | 59 |
| DS | - | - | - | - | 79 | - | 97 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| ECA | - | - | - | - | - | - | 79 | - | - | - | - | - | - | - | 72 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| ECCE | - | - | - | - | - | - | 61 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCA | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCS | - | - | - | - | - | - | 79 | - | - | 75 | - | - | - | - | - | - | 75 | - | - | - | - | - | - | - | - | - | - | - | - | 71 | - | - | - | - | - |
| FCI | 58 | 39 | 67 | 62 | 68 | 66 | 69 | 68 | 52 | 67 | 59 | 70 | 50 | 49 | 67 | 67 | - | 66 | 72 | 60 | 57 | 62 | 69 | 20 | 73 | 74 | 56 | 62 | - | 64 | 22 | 47 | 66 | - | 57 |
| FMCE | - | - | - | - | - | - | 72 | 81 | - | - | - | - | - | - | - | - | 74 | - | - | 39 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| FORT | - | - | - | - | - | - | 70 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| FTGOT | - | - | - | - | - | - | 27 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 23 | - | - |
| FVA | - | - | - | - | - | - | 76 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| GECI | - | - | - | - | - | - | 82 | - | - | - | - | - | - | - | - | - | - | - | - | 82 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| HTCE | - | - | - | - | - | - | 76 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 68 |
| IBCDC | - | - | - | - | - | - | 62 | - | - | - | - | 64 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| IBCM | - | - | - | - | - | - | 74 | - | - | - | - | 66 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LPCA | - | - | - | - | - | - | 84 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LPCI | - | - | - | - | - | - | 65 | 62 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LSCI | - | - | - | - | - | - | 73 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MBT | - | - | - | - | 41 | 44 | 44 | 40 | 27 | 53 | - | 41 | - | - | - | - | - | - | 38 | 44 | 38 | - | - | - | - | 35 | - | - | - | - | - | - | 37 | - | - |
| MCS | - | - | - | - | - | - | 58 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MUQ | - | - | - | - | - | - | 37 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MWCS1 | - | - | - | - | - | - | 79 | 76 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 35 | - | - | - |
| MWCS2 | - | - | - | - | - | - | 68 | 62 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| NGCI | 76 | - | - | - | - | - | 87 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| NGPSD | - | - | - | - | - | - | 95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| PIQL | - | - | - | - | - | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMCA5 | - | - | - | - | - | - | 74 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 67 | - | - | - | - | - | - | - | - | - |
| QMCA6 | - | - | - | - | - | - | 81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMCS | - | - | - | - | - | - | 86 | - | - | 81 | - | - | - | - | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMFPS | - | - | - | - | - | - | 60 | 49 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMS | - | - | - | - | - | - | 69 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMVI | - | - | - | - | - | - | 32 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QPCS | - | - | - | - | - | - | 71 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 40 | - | - | - |
| RAPT | - | - | - | - | - | - | 81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RCI | - | - | - | - | - | - | 75 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RFCI | - | - | - | - | - | - | 80 | - | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RKI | - | - | - | - | - | - | 70 | - | - | - | - | 67 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RRMCS | - | - | - | - | - | - | 74 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| SGCE | - | - | - | - | - | - | 60 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| SPCI | - | - | - | - | - | - | 95 | 94 | - | - | - | - | - | - | - | - | - | - | - | 86 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| STPFaSLSh | - | - | - | - | - | - | 86 | - | - | - | - | - | - | - | - | - | 70 | - | - | - | - | - | - | - | - | 72 | - | - | - | - | - | - | - | - | 59 |
| STPFaSLlo | - | - | - | - | - | - | 97 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TCE | - | - | - | - | - | - | 88 | - | - | - | - | - | - | - | - | - | - | - | - | 85 | - | - | - | - | - | 78 | - | - | - | - | - | - | - | - | 59 |
| TCS | - | - | - | - | - | - | 79 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 59 | - | - | 69 |
| TOAST | - | - | - | - | - | - | 83 | - | - | - | - | - | - | - | - | - | - | - | - | 80 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TUG-K2.6 | 48 | - | - | - | 60 | - | - | - | - | 60 | - | 67 | 43 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TUG-K3.0-4.0 | - | - | - | - | - | 44 | 55 | 60 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 65 | - | - | - | 50 | - | - | - | 45 | 45 |
| TUV | 58 | - | - | - | - | - | 60 | 62 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
Also indicated in Table 7 is the relative performance compared to published scores. For most inventories in most nominal languages, the AI system outperforms published post-instruction student scores at the undergraduate level. It is notable that for each inventory, this performance differential is fairly consistent across nominal languages.
IV.2 Score comparison by subject area
Table 8 shows a comparison of average post-test student scores found in the literature for undergraduate-level courses (see column “%Perc” in Tables 3-6) and average AI scores. We calculate the AI scores based on the scores listed in Table 7, taking the average across all languages. With the exception of laboratory skills (LAB), the AI outperforms average undergraduate students in every subject category. Note that the average AI score would usually be even higher, if we only considered tests in English.
| Category | Avg. Student Score | Avg. AI Score | AI/Students |
| AST | 48 | 79 | 1.64 |
| EM-C | 46 | 58 | 1.25 |
| EM-F | 54 | 63 | 1.17 |
| LAB | 39 | 37 | 0.94 |
| MATH | 60 | 71 | 1.19 |
| MECH | 54 | 67 | 1.24 |
| OPT | 38 | 60 | 1.57 |
| QUANT | 50 | 65 | 1.30 |
| REAS | 58 | 75 | 1.29 |
| RELA | 71 | 75 | 1.06 |
| THERM | 60 | 77 | 1.28 |
IV.3 Performance by subject area and image usage
Table 9 shows the performance by subject area and usage of images, which allows insight into the multimodal capabilities of the AI system. Here, we calculate the performance based on all questions that are in inventories that address the subject area and, for the three columns on the right, fulfill the criteria regarding the presence and use of images. For each question, we coded if it contained or referred to an image, i.e. a visual representation (sketch, graph, diagram, etc.) and if interpreting the image was required for correctly solving the task (required image), or if the image was redundant and the task could be solved by only reading the text of the problem (unneeded image).
Overall, the AI performs best in Astronomy (AST), followed by thermodynamics (THERM), relativity (RELA), and reasoning (REAS). The inventories on Astronomy tend to focus on knowledge, while THERM, RELA, and REAS inventories tend to use exact language. The AI system is particularly weak in laboratory skills (LAB), which include strategies for data collection and analysis, Optics (OPT), mechanics (MECH), and electricity and magnetism (both circuits (EM-C) and fields (EM-F)).
The performance on questions with redundant images is similar to or better than text-only questions for all subject areas, except for LAB and RELA inventories. However, in all subject areas, the AI performs worse on questions that require image interpretation than on text-only questions. RELA, OPT, MECH, MATH, and AST exhibit especially large gaps. Furthermore, while QUANT as a whole was not among the worst performing subject areas on image-based questions, QMVI, which is entirely image-based and consists predominantly of required-image problem types, was the worst-performing of all surveys in English. In line with previous research on the topic, this suggests that visual interpretation is one of the major weaknesses of the studied AI system.
| Required | Text | Unneeded | ||
| image | only | image | ||
| Subject | %Correct | %Correct | %Correct | %Correct |
| AST | 81 | 55 | 91 | 100 |
| EM-C | 53 | 52 | 61 | 100 |
| EM-F | 57 | 52 | 73 | 69 |
| LAB | 37 | 27 | 45 | 17 |
| MATH | 69 | 52 | 90 | - |
| MECH | 59 | 44 | 84 | 80 |
| OPT | 53 | 29 | 88 | 81 |
| QUANT | 64 | 55 | 66 | 79 |
| REAS | 75 | 54 | 76 | 87 |
| RELA | 75 | 0 | 85 | 67 |
| THERM | 77 | 72 | 79 | 77 |
| Overall | 63 | 48 | 81 | 79 |
IV.4 Scores by question and language
| Answers correct in EN | ||||
| 0/3 | 1/3 | 2/3 | 3/3 | |
| Language | 19% (284) | 9% (131) | 9% (128) | 64% (955) |
| AR | 17% (16) | 8% (4) | 50% (6) | 85% (50) |
| BN | 5% (7) | 0% (2) | 11% (3) | 61% (18) |
| CA | 14% (7) | 0% (2) | 78% (3) | 93% (18) |
| CS | 17% (8) | 8% (4) | 76% (7) | 88% (33) |
| DE | 14% (26) | 36% (15) | 41% (9) | 87% (46) |
| EL | 14% (36) | 43% (17) | 52% (9) | 83% (49) |
| ES | 18% (102) | 32% (42) | 56% (45) | 88% (227) |
| FA | 20% (18) | 33% (6) | 28% (6) | 59% (26) |
| FI | 24% (33) | 48% (22) | 48% (11) | 90% (85) |
| FIL | 10% (7) | 17% (2) | 44% (3) | 85% (18) |
| FR | 23% (42) | 42% (12) | 43% (21) | 89% (86) |
| HE | 10% (7) | 17% (2) | 67% (3) | 67% (18) |
| HI | 14% (7) | 0% (2) | 11% (3) | 74% (18) |
| HR | 3% (11) | 33% (6) | 67% (4) | 92% (42) |
| HU | 19% (7) | 17% (2) | 56% (3) | 93% (18) |
| ID | 25% (29) | 48% (16) | 44% (15) | 84% (107) |
| IS | 14% (7) | 0% (2) | 67% (3) | 93% (18) |
| IT | 17% (18) | 22% (6) | 56% (6) | 92% (26) |
| JA | 16% (47) | 31% (14) | 35% (17) | 83% (163) |
| MS | 9% (29) | 37% (10) | 52% (9) | 82% (40) |
| NL | 14% (7) | 0% (2) | 56% (3) | 89% (18) |
| NO | 10% (7) | 33% (2) | 56% (3) | 98% (18) |
| PA | 10% (7) | 0% (2) | 56% (3) | 20% (18) |
| PL | 29% (7) | 17% (2) | 67% (3) | 98% (18) |
| PT | 26% (44) | 37% (20) | 48% (14) | 84% (125) |
| RU | 5% (7) | 17% (2) | 67% (3) | 78% (18) |
| SK | 14% (7) | 17% (2) | 67% (3) | 85% (18) |
| SR | 67% (1) | 33% (3) | 0% (1) | 70% (19) |
| SV | 18% (49) | 36% (29) | 44% (18) | 84% (122) |
| TA | 5% (7) | 50% (2) | 22% (3) | 26% (18) |
| TH | 17% (21) | 25% (8) | 24% (7) | 59% (76) |
| TR | 15% (41) | 13% (13) | 45% (11) | 84% (31) |
| UK | 17% (10) | 0% (2) | 33% (1) | 74% (13) |
| ZH | 21% (46) | 43% (24) | 40% (15) | 73% (153) |
| Overall | ||||
| non-EN | 18% (725) | 33% (301) | 48% (274) | 82% (1771) |
| Overall | ||||
| non-EN | ||||
| Text only | 16% (51) | 52% (18) | 65% (51) | 84% (759) |
Table 10 can be used to assess if the AI struggles with the same questions across different languages. One indicator for “struggle” is if the answers for the three independent attempts agree on the correct choice: how consistent is the AI system in its correct answer choices? The by far largest number of inventories was available in English (all except for the TUG-K2.6) and we referenced the performance to questions given in English. There were 955 individual questions across the inventories for which the AI response was correct for all three responses in English. For 128 questions the AI had two out of the three responses correct. There were 131 questions with one of three correct responses and 284 with all three answers incorrect.
Table 10 shows that questions that are difficult in English tend to remain difficult in the other languages and vice versa (with the earlier mentioned exception of Punjabi and Tamil). For example, among the 102 questions for which all three English responses were incorrect, only 18% of responses in the second most common language, Spanish (ES), were correct. This increases to 32% for the 42 questions with 1/3 correct in English, 56% for the 45 questions with 2/3 correct in English, and 88% for the 227 questions where all three English responses were correct. Similar trends can be seen for the other languages and the percentages for all non-English language questions went from 18% (0/3 in EN), to 33% (1/3 in EN), to 48% (2/3 in EN), and 82% (3/3 in EN).
As discussed earlier, part of this could have been due to lower performances for questions that contain images that pose a challenge for the AI invariably of the language. However, the same trend holds if you look at the “text only” questions. While the overall performance increased (see Table 9), the correlation between questions being answered correctly in English and other languages remains. Correct responses for non-English questions went from 16% (0/3 in EN), to 52% (1/3 in EN), to 65% (2/3 in EN), and 84% (3/3 in EN).
| Incorrect | different | |||
| answers | or same | All | EN | Non-EN |
| 3 | 3 different | 8% (98) | 6% (18) | 9% (80) |
| 3 | 3 the same | 53% (617) | 61% (173) | 50% (444) |
| 3 | 2 the same | 38% (448) | 33% (93) | 40% (355) |
| 2 | 2 different | 34% (197) | 28% (37) | 36% (160) |
| 2 | 2 the same | 66% (376) | 72% (94) | 64% (282) |
Well-designed assessment instruments for teaching have multiple-choice distractors that probe particular student misconceptions. Given that we drew our questions from the PhysPort library with many instruments having been designed carefully, we would expect that for our data set students would gravitate to particular incorrect answers. As such, it is interesting to consider if the AI was gravitating toward specific incorrect answers. Table 11 shows the questions that had two or three incorrect answers for the set of the three AI responses. It then shows if the selected multiple-choice items were the same or different. The percentages for the full data set as well as the data set split into English and non-English responses show that the AI frequently picked the same incorrect multiple-choice items. While not all of the assessments analyzed in this study had five possible choices, the majority did, and we can use this as a reference. If the probability of picking a particular incorrect answer was random we would get a 75% probability of having different answers for two incorrect answers and a 25% chance of having the same answer. For our data this was flipped, 66% of all, 72% of English, and 64% of non-English questions had the same two incorrect choices. The same can be seen for cases where all three answers were incorrect. If the likelihood of a particular incorrect response was random, in 6.25% of cases all three would be the same, in 37.5% of cases all three would be different, and in 56.25% of cases two responses would be the same. As shown in Table 11, in reality, the AI responses are far from random and gravitate toward specific incorrect answers (e.g. 53%, corresponding to 618 questions, instead of 6.25% had all three incorrect responses being the same).
IV.5 Scores by language and language switching
Out of 36 tests that were translated, English was the best performing language in 27 (75%) of cases. Table 12 shows the relative performance by language, compared to English. Overall performance does depend on the language of the questions (nominal language). It is notable that for Catalan, Finnish, French, Italian, Norwegian, and Polish, the performance is comparable, with a number of languages close behind. On the other hand, for nominal languages such as Farsi, Hebrew, Hindi, and most notably Punjabi, the AI exhibits much lower performance.
| AR | BN | CA | CS | DE | EL | ES | FA | FI | FIL | FR | HE | HI | HR | HU | ID | IS | IT | JA | MS | NL | NO | PA | PL | PT | RU | SK | SR | SV | TA | TH | TR | UK | ZH | |
| ADT | - | - | - | - | - | - | 1.04 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 1.06 | - | - | - | - | - |
| BEMA | - | - | - | - | - | - | 0.75 | - | - | - | - | - | - | - | - | - | - | - | 0.80 | - | - | - | - | - | 0.87 | - | - | - | 0.77 | - | - | - | - | 0.67 |
| CCI | - | - | - | 0.95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CDPA | - | - | - | - | - | - | 1.20 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| CSEM | - | - | - | - | - | - | 0.98 | - | - | - | - | - | - | - | - | 0.69 | - | - | - | 0.98 | - | - | - | - | - | - | - | - | 0.83 | - | - | - | - | - |
| CTSR | - | - | - | - | - | - | 0.85 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.73 | 0.89 | - | - | - | - | - |
| DIRECT | - | - | - | - | 0.91 | 1.07 | 0.67 | - | 1.05 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 1.00 | - | - | - | - | 1.19 |
| DS | - | - | - | - | 0.81 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| ECA | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.91 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCA | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 1.00 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| EMCS | - | - | - | - | - | - | - | - | 0.95 | - | - | - | - | - | - | 0.95 | - | - | - | - | - | - | - | - | - | - | - | - | 0.90 | - | - | - | - | - |
| FCI | 0.84 | 0.56 | 0.97 | 0.90 | 0.98 | 0.95 | 0.98 | 0.76 | 0.97 | 0.85 | 1.02 | 0.73 | 0.71 | 0.97 | 0.97 | - | 0.95 | 1.05 | 0.87 | 0.82 | 0.90 | 1.00 | 0.29 | 1.06 | 1.08 | 0.81 | 0.90 | - | 0.94 | 0.32 | 0.68 | 0.95 | - | 0.82 |
| FMCE | - | - | - | - | - | - | 1.13 | - | - | - | - | - | - | - | - | 1.04 | - | - | 0.54 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| FTGOT | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.87 | - | - |
| GECI | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 1.00 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| HTCE | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.89 |
| IBCDC | - | - | - | - | - | - | - | - | - | - | 1.03 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| IBCM | - | - | - | - | - | - | - | - | - | - | 0.89 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| LPCI | - | - | - | - | - | - | 0.95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| MBT | - | - | - | - | 0.94 | 1.00 | 0.91 | 0.62 | 1.21 | - | 0.94 | - | - | - | - | - | - | 0.88 | 1.00 | 0.88 | - | - | - | - | 0.79 | - | - | - | - | - | - | 0.85 | - | - |
| MWCS1 | - | - | - | - | - | - | 0.96 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.44 | - | - | - |
| MWCS2 | - | - | - | - | - | - | 0.91 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| NGCI | 0.87 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMCA5 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.90 | - | - | - | - | - | - | - | - | - |
| QMCS | - | - | - | - | - | - | - | - | 0.94 | - | - | - | - | - | - | - | - | - | 0.90 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QMFPS | - | - | - | - | - | - | 0.82 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| QPCS | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.57 | - | - | - |
| RFCI | - | - | - | - | - | - | - | - | 1.00 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| RKI | - | - | - | - | - | - | - | - | - | - | 0.95 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| SPCI | - | - | - | - | - | - | 0.98 | - | - | - | - | - | - | - | - | - | - | - | 0.90 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| STPFaSLSh | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.81 | - | - | - | - | - | - | - | - | 0.84 | - | - | - | - | - | - | - | - | 0.68 |
| TCE | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.96 | - | - | - | - | - | 0.88 | - | - | - | - | - | - | - | - | 0.67 |
| TCS | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.75 | - | - | 0.87 |
| TOAST | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | 0.97 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| TUG-K | 0.86 | - | - | - | 1.09 | 0.79 | 1.09 | - | 1.09 | - | 1.21 | 0.78 | - | - | - | - | - | - | - | - | - | - | - | - | 1.19 | - | - | - | 0.91 | - | - | - | 0.81 | 0.81 |
| TUV | 0.97 | - | - | - | - | - | 1.03 | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - | - |
| Average | 0.89 | 0.56 | 0.97 | 0.92 | 0.95 | 0.95 | 0.95 | 0.69 | 1.03 | 0.85 | 1.01 | 0.75 | 0.71 | 0.94 | 0.97 | 0.90 | 0.95 | 0.97 | 0.88 | 0.90 | 0.90 | 1.00 | 0.29 | 1.06 | 0.94 | 0.81 | 0.90 | 0.73 | 0.91 | 0.32 | 0.61 | 0.89 | 0.81 | 0.83 |
It is unclear in which language the AI system internally processes the problems, however, the language of the output may be a good indicator, which exhibits code-switching. As seen in Table 14, for some languages, the AI performs better when the output is either entirely in English or mixed with English. Languages like Bengali, Persian, and Punjabi have a marked increase in accuracy when the AI switches to English, implying difficulties in processing the original language of the inventories. Meanwhile, languages like English, Spanish, and French demonstrate more balanced performance, with nominal language usage still resulting in high correctness scores. If the performance of the AI system for a particular question in English is used as a measure of difficulty, Table 13 indicates that the AI stays in the nominal language more often for easier questions. For all non-English questions combined, it does so for 24% of the questions that are deemed easier in English (3/3 correct answers in English). While not a strong signal, it is markedly more than 15-17% for the questions where at least one response in English was incorrect.
| Answers correct in EN | ||||
| Language switch | 0/3 | 1/3 | 2/3 | 3/3 |
| Switch to English | 69% | 69% | 69% | 63% |
| Mixed Language | 14% | 17% | 15% | 12% |
| Nominal Language | 17% | 15% | 17% | 24% |
Clustering emerges if we consider how often different languages remain in their nominal language versus switching to English or mixing languages; Fig. 7 shows a heat map of the cosine-similarity between the tuples (%nominal language, %switching to English, %mixed language) for each language. Spanish and Portuguese form a cluster that in the majority of cases stays within the nominal languages; the two languages are also part of a larger cluster of commonly spoken European languages, namely German, Italian, Dutch, and French. The next cluster includes several Scandinavian and Eastern European languages. Finally, a large cluster of mostly Asian and Arabic languages emerges, which with the notable exceptions of Polish and Icelandic uses non-Latin script. Thus, from a linguistic standpoint, we can group languages based on their dependence on language switching: those that remain robust in their native form, those that need English to boost correctness, and those that benefit from mixing strategies. This language-based clustering mirrors the subject-area clustering by revealing underlying differences in how the AI system processes and understands content under different linguistic and visual conditions.
| Nominal language | Switch to English | Mixed languages | |||||
| Language | %Correct | % | %Correct | % | %Correct | % | %Correct |
| AR | 60 | 6 | 89 | 92 | 58 | 2 | 83 |
| BN | 39 | 2 | 100 | 98 | 38 | 0 | |
| CA | 67 | 19 | 94 | 60 | 57 | 21 | 68 |
| CS | 69 | 3 | 100 | 58 | 54 | 40 | 89 |
| DE | 56 | 19 | 70 | 40 | 49 | 41 | 56 |
| EL | 52 | 8 | 63 | 80 | 50 | 11 | 61 |
| EN | 72 | 100 | 72 | - | - | - | - |
| ES | 62 | 59 | 63 | 33 | 62 | 8 | 52 |
| FA | 40 | 1 | 0 | 98 | 40 | 1 | 100 |
| FI | 66 | 4 | 86 | 60 | 65 | 36 | 64 |
| FIL | 59 | 0 | 100 | 59 | 0 | ||
| FR | 63 | 37 | 74 | 43 | 55 | 20 | 60 |
| HE | 47 | 1 | 100 | 97 | 46 | 2 | 67 |
| HI | 49 | 9 | 75 | 90 | 47 | 1 | 0 |
| HR | 69 | 11 | 81 | 64 | 69 | 25 | 66 |
| HU | 67 | 9 | 100 | 70 | 65 | 21 | 58 |
| ID | 67 | 26 | 71 | 59 | 67 | 15 | 60 |
| IS | 66 | 0 | 100 | 66 | 0 | ||
| IT | 57 | 37 | 71 | 53 | 48 | 10 | 47 |
| JA | 63 | 8 | 87 | 83 | 60 | 8 | 72 |
| MS | 50 | 0 | 100 | 50 | 0 | ||
| NL | 62 | 34 | 68 | 39 | 46 | 27 | 79 |
| NO | 69 | 13 | 75 | 71 | 62 | 16 | 93 |
| PA | 20 | 2 | 100 | 93 | 19 | 4 | 0 |
| PL | 73 | 0 | 100 | 73 | 0 | ||
| PT | 64 | 56 | 73 | 34 | 50 | 10 | 63 |
| RU | 56 | 8 | 57 | 84 | 55 | 8 | 57 |
| SK | 62 | 0 | 76 | 62 | 24 | 64 | |
| SR | 62 | 0 | 100 | 62 | 0 | ||
| SV | 60 | 13 | 65 | 56 | 59 | 31 | 58 |
| TA | 22 | 1 | 100 | 87 | 19 | 12 | 36 |
| TH | 47 | 7 | 67 | 86 | 44 | 7 | 55 |
| TR | 40 | 18 | 51 | 64 | 39 | 18 | 35 |
| UK | 45 | 3 | 0 | 79 | 45 | 18 | 50 |
| ZH | 58 | 2 | 57 | 97 | 58 | 1 | 20 |
| Overall | 63 | 46 | 71 | 45 | 55 | 9 | 60 |
V Discussion
The results presented here highlight several noteworthy patterns in the AI’s performance across different physics inventories, languages, and question formats. Unlike many previous studies where inventories were provided as text-only materials, this study supplied the AI with a screenshot of the question the way a student would see it. This richer, more realistic presentation included the actual layout, images, and language-specific formatting, effectively increasing the complexity of the input. The AI’s decision-making process—whether to remain in the nominal language, switch entirely to English, or produce a mixed-language response—was an emergent behavior rather than an externally imposed constraint. The prompt itself was given in English, but the question content was often in another language, with images embedded in the screenshot. Thus, the AI’s language choices and solution correctness provide insights into its internal reasoning and capability to handle diverse formats.
Across the majority of inventories and subject areas (with laboratory skills being the exception), the AI achieves higher scores and performs better than the average undergraduate student post-instruction, based on published results. It needs to be emphasized that the instruments are research-based and thus carefully designed with regard to psychometric properties, which is reflected in their score distributions:
The test items aim for medium difficulty, resulting in average scores around 50%.
The test items are designed for high discrimination, resulting in broad score distributions.
Thus, the finding that the AI performs better than the average published results, in broad strokes, means that it performs better than 50% on these instruments. It also means that there will almost always be students who outperform the AI. Whenever a study included results for graduate-student or post-doctoral populations, these clearly outperformed the AI.
Looking at the presence and function of visual representations in questions, we can see a clear pattern. The AI performed better on text-only questions and questions with redundant visual representations, compared to questions where visual interpretation of images was necessary for solving the task correctly. This was true for each individual subject area, as well as across the entire dataset. The average difference in performance between required image and text-only questions was 33 percentage points, with some subject areas exhibiting much higher gaps (see Table 9). This corroborates previous findings on the topic, which suggest that the AI’s vision abilities present a major weakness and hamper its ability to engage with physics tasks that require the interpretation in graphical or picture formats commonly found in physics [24, 31]. For at least the present time, instructors who want to discourage students from using AI to do their homework, including required information in some kind of graphical or picture (or even video) format is good advice.
On the language side, the results show a complex interplay between linguistic capability and the AI’s self-directed switching behavior. With the exception of Spanish and Portuguese, the AI switches away from the nominal language in the majority of cases. This outcome likely points to limitations in the AI’s mastery of certain languages, where staying in the original language of the prompt results in lower accuracy. Conversely, in languages such as German, French, Spanish, and Portuguese, the AI performs better on questions where it stays in the nominal language, which may have at least two possible explanations:
The AI is more adept at these languages or that their training data better captures the relevant scientific vocabulary and discourse, or
The AI switches languages for the problems where it already “struggles,” and then performs worse on those than on the ones where straightforward inferences can be found in the nominal language.
The fact that the AI autonomously decided to switch or mix languages further suggests that it was “aware,” at least statistically, of its varying strengths and weaknesses across linguistic contexts.
Our results show that questions that are difficult in English are likely difficult in other languages as well. Either the AI switched to English in its reasoning or, when staying in the nominal language, as frequently seen in Spanish, similar questions were difficult (or easy) for the AI even though the training data sets in the different languages were not the same. Like students, the AI gravitated toward particular incorrect solutions. How the level of difficulty for a specific question type is different for students versus AI is intriguing. When going astray, do students and AI gravitate toward the same incorrect answers? Are the preferred incorrect AI selections a reflection of issues in the training data or just more challenging for the logic built into GPT-4o? While this study does not answer these questions, it provides a stepping stone for future research. Our study indicates we can identify questions that are consistently challenging for AI across languages. Designers and instructors could specifically look at assessment designs that are challenging for AIs. The most obvious is the inclusion of images but this work clearly shows text-based questions can also pose challenges across languages for a multimodal AI system.
Overall, the study’s methodology – providing authentic screenshots rather than merely text and having the AI itself choose its linguistic mode – brings us closer to understanding how such systems would behave in real-world educational settings. We see that certain subject matters are intrinsically more challenging, that language understanding varies widely, and that the AI’s own adaptive language strategies can partially compensate for these differences. Nonetheless, large gaps and outliers remain, pointing to areas where improvements in training data, prompt design, or model architecture are needed. Future work should focus on exploring how to guide the system’s language-switching decisions more effectively, refining the visual understanding components, and ensuring more balanced coverage across different languages and conceptual domains.
VI Limitations
This study is decidedly exploratory and empirical. The preparation of question images was done manually, and some manual cleanup of the data was required. Given more than 3,600 images and the random oddities occurring in over 14,000 solutions generated by a probabilistic system, clerical errors cannot be excluded. Additionally, each screenshot of test items was iterated only three times: given the stochastic nature of LLM outputs, this introduces variability that prevents the results from being generalizable. The study also did not consider the quality of the concept test translations, and lower scores may be due to incorrect or confusing translations.
Performance might be dependent on prompts, and the ones in Figs. 3 and 4 may not be the best choices.
It should be noted that since we did not score the correctness of the physics reasoning in the AI answer, a key stumbling block for the AI can be the processing of graphical information unrelated to a physics concept or language. For example, for Question 7 in the FCI survey, the multiple-choice options are embedded in a graphic showing a ball swung in a circular path and not listed separately in the text. Based on the problem description in the text, the AI frequently gave physically reasonable answers but for the 32 languages and 96 total answers, it selected the correct choice that was embedded as a letter or number in the graphic just once. For no obvious reason, it chose instead one particular incorrect answer an eyebrow-raising 88 times. This confirms previous studies that have shown that the graphical layout of images and the spatial arrangement of answer options can play an important role in the AI’s selection of answers [31].
Finally, human post-instruction scores were gathered on a best-effort base, which may also introduce variability into the comparisons.
VII Conclusion
The results of this study underscore the complexity and variability inherent in using multimodal large language models for physics assessment tasks across multiple languages and formats. By presenting the LLM with authentic screenshots of physics concept inventories, we moved beyond text-only evaluations and closer to the conditions encountered by students in real educational settings. The AI’s own decision to switch, mix, or maintain the nominal language of the problem allowed us to gain insights into its internal reasoning and its adaptive strategies when confronted with linguistic or conceptual challenges.
While the AI’s performance depends on the subject area, the reasons for this are not entirely clear. Possible explanations include different levels of representation in the training data and its varying quality across the subject areas, or potential differences in the inherent difficulty of tasks in the assessments covering different subject areas. However, we have found that one variable clearly influencing AI’s performance across all subject areas is the presence of non-redundant visual representations. This suggests that the AI’s vision abilities still present a major weakness in its engagement with physics problems.
The marked differences in performance across languages and language-switching strategies highlight that LLMs are not equally competent in all tongues. Languages with less representation in training data often see dramatic performance gains when the AI switches to English. Meanwhile, languages better supported by training data and consistent physics terminology show more stable results. This observation not only illuminates the AI’s uneven linguistic capabilities but also the influence of translation quality and cultural-linguistic nuances on physics problem-solving accuracy.
The results also show that providing the AI with more authentic, image-based input adds another layer of complexity. The need for the AI system to interpret visual information and integrate it with textual cues can strain its reasoning process, leading to uneven performance when images are integral to the question. This finding is critical when considering the use of LLMs as learning aids or assessment tools: while their linguistic capabilities are evident, their visual and multimodal reasoning abilities remain limited and domain-dependent.
Based on published scores of post-instruction deployment of the inventories, we found that AI system outperforms average undergraduate students in all subject areas but laboratory skills. We were not able to locate sufficient language-specific data for students, but found that the AI-system’s outperforming of undergraduate students is largely independent of the nominal language of the inventory.
In sum, this study demonstrates that while the studied AI system is capable and adaptive in many respects, it still exhibits significant variations in performance depending on the language, conceptual domain, and the presence of visual information. The work points toward future improvements in training data diversity, model fine-tuning, and prompt engineering to enhance both linguistic and conceptual understanding, as well as visual interpretation of multiple representations. It also highlights the need for careful consideration when implementing such AI systems in educational contexts, ensuring that their use is both equitable and aligned with pedagogical goals.
Data Availability
To be determined.
Acknowledgements.
We would like to thank Sam McKagan for all of her support.References
- Vaswani et al. [2017] A. Vaswani, N. Shazeer, N. Parmar, J. Uszkoreit, L. Jones, A. N. Gomez, Ł. Kaiser, and I. Polosukhin, Attention is all you need, Advances in neural information processing systems 30 (2017).
- Kung et al. [2022] T. H. Kung, M. Cheatham, A. Medinilla, ChatGPT, C. Sillos, L. De Leon, C. Elepano, M. Madriaga, R. Aggabao, G. Diaz-Candido, et al., Performance of chatgpt on usmle: Potential for ai-assisted medical education using large language models, medRxiv , 2022 (2022).
- Samantha Murphy Kelly [2023] Samantha Murphy Kelly, ChatGPT passes exams from law and business schools, https://edition.cnn.com/2023/01/26/tech/chatgpt-passes-exams/index.html (accessed January 2023).
- OpenAI [2024a] OpenAI, ChatGPT, https://chat.openai.com/ (accessed April 2024a).
- Turing [1950] A. M. Turing, Computing machinery and intelligence, Mind , 433 (1950).
- Jones and Bergen [2024] C. R. Jones and B. K. Bergen, People cannot distinguish gpt-4 from a human in a turing test, arXiv preprint arXiv:2405.08007 (2024).
- Achiam et al. [2023] J. Achiam, S. Adler, S. Agarwal, L. Ahmad, I. Akkaya, F. L. Aleman, D. Almeida, J. Altenschmidt, S. Altman, S. Anadkat, et al., GPT-4 technical report, arXiv preprint arXiv:2303.08774 (2023).
- OpenAI [2024b] OpenAI, ChatGPT, https://openai.com/research/gpt-4 (accessed April 2024b).
- Kortemeyer [2023a] G. Kortemeyer, Could an artificial-intelligence agent pass an introductory physics course?, Phys. Rev. Phys. Educ. Res. 19, 010132 (2023a).
- Polverini and Gregorcic [2024a] G. Polverini and B. Gregorcic, How understanding large language models can inform the use of chatgpt in physics education, European Journal of Physics 45, 025701 (2024a).
- Kortemeyer and Bauer [2024] G. Kortemeyer and W. Bauer, Cheat sites and artificial intelligence usage in online introductory physics courses: What is the extent and what effect does it have on assessments?, Phys. Rev. Phys. Educ. Res. 20, 010145 (2024).
- Yeadon and Hardy [2024] W. Yeadon and T. Hardy, The impact of AI in physics education: a comprehensive review from GCSE to university levels, Physics Education 59, 025010 (2024).
- Pimbblet and Morrell [2024] K. A. Pimbblet and L. J. Morrell, Can chatgpt pass a physics degree? making a case for reformation of assessment of undergraduate degrees, European Journal of Physics 46, 015702 (2024).
- Sperling and Lincoln [2024] A. Sperling and J. Lincoln, Artificial intelligence and high school physics, The Physics Teacher 62, 314 (2024).
- Bitzenbauer [2023] P. Bitzenbauer, Chatgpt in physics education: A pilot study on easy-to-implement activities, Contemporary Educational Technology 15, ep430 (2023).
- Küchemann et al. [2023] S. Küchemann, S. Steinert, N. Revenga, M. Schweinberger, Y. Dinc, K. E. Avila, and J. Kuhn, Can ChatGPT support prospective teachers in physics task development?, Phys. Rev. Phys. Educ. Res. 19, 020128 (2023).
- Kortemeyer [2023b] G. Kortemeyer, Using artificial-intelligence tools to make LaTeX content accessible to blind readers, TUGboat 44, 390 (2023b).
- Gregorcic et al. [2024] B. Gregorcic, G. Polverini, and A. Sarlah, Chatgpt as a tool for honing teachers’ socratic dialogue skills, Physics Education 59, 045005 (2024).
- Crawford et al. [2023] J. Crawford, M. Cowling, and K.-A. Allen, Leadership is needed for ethical chatgpt: character, assessment, and learning using artificial intelligence (ai), Journal of University Teaching and Learning Practice 20 (2023).
- Vasconcelos and Dos Santos [2023] M. A. R. Vasconcelos and R. P. Dos Santos, Enhancing stem learning with chatgpt and bing chat as objects to think with: a case study, Eurasia Journal of Mathematics, Science and Technology Education 19, em2296 (2023).
- Dahlkemper et al. [2023] M. N. Dahlkemper, S. Z. Lahme, and P. Klein, How do physics students evaluate artificial intelligence responses on comprehension questions? a study on the perceived scientific accuracy and linguistic quality of ChatGPT, Phys. Rev. Phys. Educ. Res. 19, 010142 (2023).
- Ding et al. [2023] L. Ding, T. Li, S. Jiang, and A. Gapud, Students’ perceptions of using ChatGPT in a physics class as a virtual tutor, International Journal of Educational Technology in Higher Education 20, 63 (2023).
- West [2023] C. G. West, AI and the FCI: Can ChatGPT project an understanding of introductory physics? (2023), arXiv:2303.01067 [physics.ed-ph] .
- Polverini and Gregorcic [2024b] G. Polverini and B. Gregorcic, Evaluating vision-capable chatbots in interpreting kinematics graphs: a comparative study of free and subscription-based models, in Frontiers in Education, Vol. 9 (Frontiers Media SA, 2024) p. 1452414.
- OpenAI [2024c] OpenAI, Hello GPT-4o, https://openai.com/index/hello-gpt-4o/ (accessed June 2024c).
- Hestenes et al. [1992] D. Hestenes, M. Wells, and G. Swackhamer, Force Concept Inventory, The Physics Teacher 30, 141 (1992), https://doi.org/10.1119/1.2343497 .
- Beichner [1994] R. J. Beichner, Testing student interpretation of kinematics graphs, American journal of Physics 62, 750 (1994).
- Ding et al. [2006] L. Ding, R. Chabay, B. Sherwood, and R. Beichner, Evaluating an electricity and magnetism assessment tool: Brief electricity and magnetism assessment, Phys. Rev. ST Phys. Educ. Res. 2, 010105 (2006).
- Aldazharova et al. [2024] S. Aldazharova, G. Issayeva, S. Maxutov, and N. Balta, Assessing AI’s problem solving in physics: Analyzing reasoning, false positives and negatives through the force concept inventory, Contemporary Educational Technology 16, ep538 (2024).
- Polverini and Gregorcic [2024c] G. Polverini and B. Gregorcic, Performance of chatgpt on the test of understanding graphs in kinematics, Phys. Rev. Phys. Educ. Res. 20, 010109 (2024c).
- Polverini et al. [2024] G. Polverini, J. Melin, E. Önerud, and B. Gregorcic, Performance of chatgpt on tasks involving physics visual representations: the case of the brief electricity and magnetism assessment, arXiv preprint arXiv:2412.10019 (2024).
- Henderson [2002] C. Henderson, Common concerns about the force concept inventory, The Physics Teacher 40, 542 (2002).
- Gregorcic and Pendrill [2023] B. Gregorcic and A.-M. Pendrill, ChatGPT and the frustrated Socrates, Physics Education 58, 035021 (2023).
- Madsen et al. [2017] A. Madsen, S. B. McKagan, and E. C. Sayre, Best practices for administering concept inventories, The Physics Teacher 55, 530 (2017).
- Hake [1998] R. R. Hake, Interactive-engagement versus traditional methods: A six-thousand-student survey of mechanics test data for introductory physics courses, American journal of Physics 66, 64 (1998).
- Sands et al. [2018] D. Sands, M. Parker, H. Hedgeland, S. Jordan, and R. Galloway, Using concept inventories to measure understanding, Higher Education Pedagogies 3, 173 (2018).
- Laverty and Caballero [2018] J. T. Laverty and M. D. Caballero, Analysis of the most common concept inventories in physics: What are we assessing?, Physical Review Physics Education Research 14, 010123 (2018).
- Stoen et al. [2020] S. M. Stoen, M. A. McDaniel, R. F. Frey, K. M. Hynes, and M. J. Cahill, Force concept inventory: More than just conceptual understanding, Physical Review Physics Education Research 16, 010105 (2020).
- OpenAI [2024d] OpenAI, GPT-4, https://openai.com/index/gpt-4-research/ (accessed December 2024d).
- Christiano et al. [2017] P. F. Christiano, J. Leike, T. Brown, M. Martic, S. Legg, and D. Amodei, Deep reinforcement learning from human preferences, Advances in neural information processing systems 30 (2017).
- Deepseek [2024] Deepseek, https://www.deepseek.com/ (accessed December 2024).
- Alibaba Cloud [2024] Alibaba Cloud, Qwen, https://qwen-ai.com/ (accessed December 2024).
- AI [2025] S. AI, Swiss AI Initiative, https://www.swiss-ai.org/ (retrieved January 2025).
- Nicholas and Bhatia [2023] G. Nicholas and A. Bhatia, Lost in translation: large language models in non-english content analysis, arXiv preprint arXiv:2306.07377 (2023).
- Cohere for AI [2024] Cohere for AI, The AI language gap, https://cohere.com/research/papers/the-ai-language-gap.pdf (accessed December 2024).
- Feng et al. [2024] S. Feng, W. Shi, Y. Wang, W. Ding, O. Ahia, S. S. Li, V. Balachandran, S. Sitaram, and Y. Tsvetkov, Teaching LLMs to abstain across languages via multilingual feedback, arXiv preprint arXiv:2406.15948 (2024).
- Cho [2024] N. Cho, An investigation of using Spark generative AI in solving physics concept inventories in english and chinese: Performance and issues, Discover Artificial Intelligence 4, 1 (2024).
- Martin-Jones [1995] M. Martin-Jones, Code-switching in the classroom: Two decades of research, One speaker, two languages: Cross-disciplinary perspectives on code-switching , 90 (1995).
- Myers-Scotton [2017] C. Myers-Scotton, Code-switching, The handbook of sociolinguistics , 217 (2017).
- Andersson and Rusanganwa [2011] I. Andersson and J. Rusanganwa, Language and space in a multilingual undergraduate physics classroom in rwanda, International Journal of Bilingual Education and Bilingualism 14, 751 (2011).
- Borg [2024] N. A. Borg, The role of language in physics education in bilingual and multilingual classrooms, Malta Journal of Education 5, 6 (2024).
- American Association of Physics Teachers [2017] American Association of Physics Teachers, Physport, https://www.physport.org (2017), [retrieved November 2024].
- McKagan et al. [2020] S. B. McKagan, L. E. Strubbe, L. J. Barbato, B. A. Mason, A. M. Madsen, and E. C. Sayre, PhysPort use and growth: Supporting physics teaching with research-based resources since 2011, The Physics Teacher 58, 465 (2020).
- International Organization for Standardization [2024] International Organization for Standardization, ISO 639:2023 code for individual languages and language groups, https://www.iso.org/obp/ui/en/#iso:std:iso:639:ed-2:v1:en (accessed November 2024).
- Von Korff et al. [2016] J. Von Korff, B. Archibeque, K. A. Gomez, T. Heckendorf, S. B. McKagan, E. C. Sayre, E. W. Schenk, C. Shepherd, and L. Sorell, Secondary analysis of teaching methods in introductory physics: A 50 k-student study, American Journal of physics 84, 969 (2016).
- Hufnagel [2002] B. Hufnagel, Development of the astronomy diagnostic test, Astronomy Education Review 1, 47 (2002).
- Brogt et al. [2007] E. Brogt, D. Sabers, E. E. Prather, G. L. Deming, B. Hufnagel, and T. F. Slater, Analysis of the astronomy diagnostic test, Astronomy Education Review 6, 25 (2007).
- Koca et al. [2024] N. O. Koca, N. alhuda Al Saqri, H. Al Hamrashdi, and N. Al Kindi, Evaluating the students’ learning on the electricity and magnetism using a conceptual survey bema, Physics Education 60, 015022 (2024).
- Pollock [2008] S. J. Pollock, Comparing student learning with multiple research-based conceptual surveys: Csem and bema., in AIP Conference Proceedings, Vol. 1064 (American Institute of Physics, 2008) pp. 171–174.
- Epstein [2006] J. Epstein, The calculus concept inventory, National STEM Assessment, Washington, DC , 60 (2006).
- Maciejewski [2016] W. Maciejewski, Flipping the calculus classroom: an evaluative study, Teaching Mathematics and its Applications: An International Journal of the IMA 35, 187 (2016).
- Day and Bonn [2011] J. Day and D. Bonn, Development of the concise data processing assessment, Phys. Rev. ST Phys. Educ. Res. 7, 010114 (2011).
- Maloney et al. [2001] D. P. Maloney, T. L. O’Kuma, C. J. Hieggelke, and A. Van Heuvelen, Surveying students’ conceptual knowledge of electricity and magnetism, American Journal of Physics 69, S12 (2001).
- Tapping et al. [2019] R. Tapping, G. Lepage, and N. Holmes, Visualizing patterns in csem responses to assess student conceptual understanding, in 2018 Physics Education Research Conference (PERC) (2019) pp. 419–422.
- Lawson [1978] A. E. Lawson, The development and validation of a classroom test of formal reasoning., Journal of Research in Science Teaching (1978).
- Moore and Rubbo [2012] J. C. Moore and L. J. Rubbo, Scientific reasoning abilities of nonscience majors in physics-based courses, Physical Review Special Topics—Physics Education Research 8, 010106 (2012).
- Engelhardt and Beichner [2004] P. V. Engelhardt and R. J. Beichner, Students’ understanding of direct current resistive electrical circuits, American journal of physics 72, 98 (2004).
- Sangam and Jesiek [2012] D. Sangam and B. K. Jesiek, Conceptual understanding of resistive electric circuits among first-year engineering students, in 2012 ASEE Annual Conference & Exposition (2012) pp. 25–339.
- Yeend et al. [2001] R. Yeend, M. Loverude, and B. Gonzales, Student understanding of density: a cross-age investigation, in Physics Education Research Conference (2001).
- Zenger and Bitzenbauer [2022] T. Zenger and P. Bitzenbauer, Exploring german secondary school students’ conceptual knowledge of density, Science Education International 33, 86 (2022).
- Ding et al. [2013] L. Ding, R. Chabay, and B. Sherwood, How do students in an innovative principle-based mechanics course understand energy concepts?, Journal of research in science teaching 50, 722 (2013).
- Sokoloff [1996] D. R. Sokoloff, Teaching electric circuit concepts using microcomputer-based current/voltage probes, in Microcomputer–based labs: Educational research and standards (Springer, 1996) pp. 129–146.
- Kortemeyer et al. [2019] G. Kortemeyer, D. Anderson, A. M. Desrochers, A. Hackbardt, K. Hoekstra, A. Holt, A. Iftekhar, T. Kabaker, N. Keller, Z. Korzecke, et al., Using a computer game to teach circuit concepts, European Journal of Physics 40, 055703 (2019).
- McColgan et al. [2017] M. W. McColgan, R. A. Finn, D. L. Broder, and G. E. Hassel, Assessing students’ conceptual knowledge of electricity and magnetism, Physical Review Physics Education Research 13, 020121 (2017).
- Singh and Rosengrant [2003] C. Singh and D. Rosengrant, Multiple-choice test of energy and momentum concepts, American Journal of Physics 71, 607 (2003), https://pubs.aip.org/aapt/ajp/article-pdf/71/6/607/7531054/607_1_online.pdf .
- Sahin [2010] M. Sahin, The impact of problem-based learning on engineering students’ beliefs about physics and conceptual understanding of energy and momentum, European Journal of Engineering Education 35, 519 (2010).
- Mason [2020] A. J. Mason, Learning goals and perceived irrelevance to major within life science majors in introductory physics, arXiv preprint arXiv:2012.09898 (2020).
- Kortemeyer [2009] G. Kortemeyer, Gender differences in the use of an online homework system in an introductory physics course, Phys. Rev. ST Phys. Educ. Res. 5, 010107 (2009).
- Han et al. [2015] J. Han, L. Bao, L. Chen, T. Cai, Y. Pi, S. Zhou, Y. Tu, and K. Koenig, Dividing the force concept inventory into two equivalent half-length tests, Phys. Rev. ST Phys. Educ. Res. 11, 010112 (2015).
- Thornton and Sokoloff [1998] R. K. Thornton and D. R. Sokoloff, Assessing student learning of newton’s laws: The force and motion conceptual evaluation and the evaluation of active learning laboratory and lecture curricula, American Journal of Physics 66, 338 (1998).
- Cummings et al. [1999] K. Cummings, J. Marx, R. Thornton, and D. Kuhl, Evaluating innovation in studio physics, American journal of physics 67, S38 (1999).
- Kalinowski and Willoughby [2019] S. T. Kalinowski and S. Willoughby, Development and validation of a scientific (formal) reasoning test for college students, Journal of Research in Science Teaching 56, 1269 (2019).
- Kaltakci-Gurel et al. [2017] D. Kaltakci-Gurel, A. Eryilmaz, and L. C. McDermott, Development and application of a four-tier test to assess pre-service physics teachers’ misconceptions about geometrical optics, ReseaRch in science & Technological educaTion 35, 238 (2017).
- Rosenblatt and Heckler [2011] R. Rosenblatt and A. F. Heckler, Systematic study of student understanding of the relationships between the directions of force, velocity, and acceleration in one dimension, Physical Review Special Topics - Physics Education Research 7, 020112 (2011).
- Keller [2006] J. M. Keller, Part I. development of a concept inventory addressing students’ beliefs and reasoning difficulties regarding the greenhouse effect, part II. distribution of chlorine measured by the mars odyssey gamma ray spectrometer (The University of Arizona, 2006).
- Tanahoung et al. [2006] C. Tanahoung, M. D. Sharma, I. D. Johnston, R. Chitaree, and C. Soankwan, Surveying sydney introductory physics students’ understandings of heat and temperature, in Australian Institute of Physics 17th National Congress, Brisbane, Paper No. WC0233 (2006).
- Halloun [2007] I. Halloun, Evaluation of the impact of the new physics curriculum on the conceptual profiles of secondary students. 1–25, https://www.halloun.net/wp-content/uploads/2016/10/LU-Summative-Report-10-07.pdf (2007).
- Ndihokubwayo et al. [2020] K. Ndihokubwayo, J. Uwamahoro, I. Ndayambaje, and M. Ralph, Light phenomena conceptual assessment: an inventory tool for teachers, Physics Education 55, 035009 (2020).
- Lindell and Olsen [2002] R. S. Lindell and J. P. Olsen, Developing the lunar phases concept inventory, in Proceedings of the 2002 Physics Education Research Conference (New York: PERC Publishing, 2002).
- Bardar et al. [2007] E. M. Bardar, E. E. Prather, K. Brecher, and T. F. Slater, Development and validation of the light and spectroscopy concept inventory, Astronomy Education Review 5, 103 (2007).
- Wallace et al. [2018] C. S. Wallace, T. G. Chambers, and E. E. Prather, Item response theory evaluation of the light and spectroscopy concept inventory national data set, Physical Review Physics Education Research 14, 010149 (2018).
- Hestenes and Wells [1992] D. Hestenes and M. Wells, A mechanics baseline test, The physics teacher 30, 159 (1992).
- Millán and Otranto [2021] C. P. Millán and S. Otranto, Thirty-six years of the forced concept inventory and the mechanics baseline test: is aristotle still playing hide and seek in our classrooms?, Latin-American Journal of Physics Education 15, 9 (2021).
- Antwi et al. [2011] V. Antwi, R. Hanson, A. Sam, E. Savelsbergh, and H. Eijkelhof, The impact of interactive-engagement (ie) teaching on students understanding of concepts in mechanics: The use of force concept inventory (fci) and mechanics baseline test (mbt), International Journal of Educational Planning & Administration 1, 81 (2011).
- Kádár and Tasnádi [2019] C. Kádár and P. Tasnádi, The knowledge of hungarian students in the light of the mechanics baseline test, in Journal of Physics: Conference Series, Vol. 1286 (IOP Publishing, 2019) p. 012026.
- Li and Singh [2016] J. Li and C. Singh, Developing and validating a conceptual survey to assess introductory physics students’ understanding of magnetism, European Journal of Physics 38, 025702 (2016).
- Deardorff [2001] D. L. Deardorff, Introductory physics students’ treatment of measurement uncertainty (North Carolina State University, 2001).
- Tongchai et al. [2009] A. Tongchai, M. D. Sharma, I. D. Johnston, K. Arayathanitkul, and C. Soankwan, Developing, evaluating and demonstrating the use of a conceptual survey in mechanical waves, International Journal of Science Education 31, 2437 (2009).
- Santoso et al. [2022] P. H. Santoso, E. Istiyono, and H. Haryanto, Principal component analysis and exploratory factor analysis of the mechanical waves conceptual survey, JP3I (Jurnal Pengukuran Psikologi dan Pendidikan Indonesia) 11, 209 (2022).
- Williamson et al. [2013] K. E. Williamson, S. D. Willoughby, and E. E. Prather, Development of the Newtonian gravity concept inventory, Astron. Educ. Rev. 12, 1 (2013).
- Engelhardt et al. [2018] P. V. Engelhardt, S. Robinson, E. P. Price, P. S. Smith, and F. Goldberg, Developing a conceptual assessment for a modular curriculum, in Physics Education Research Conference (2018).
- White Brahmia et al. [2021] S. White Brahmia, A. Olsho, T. I. Smith, A. Boudreaux, P. Eaton, and C. Zimmerman, Physics inventory of quantitative literacy: A tool for assessing mathematical reasoning in introductory physics, Physical Review Physics Education Research 17, 020129 (2021).
- Sadaghiani and Pollock [2015] H. R. Sadaghiani and S. J. Pollock, Quantum mechanics concept assessment: Development and validation study, Physical Review Special Topics-Physics Education Research 11, 010110 (2015).
- McKagan et al. [2010] S. McKagan, K. Perkins, and C. Wieman, Design and validation of the quantum mechanics conceptual survey, Physical Review Special Topics – Physics Education Research 6, 020121 (2010).
- Marshman and Singh [2019] E. Marshman and C. Singh, Validation and administration of a conceptual survey on the formalism and postulates of quantum mechanics, Physical Review Physics Education Research 15, 020128 (2019).
- Marshman [2015] E. M. Marshman, Improving the quantum mechanics content knowledge and pedagogical content knowledge of physics graduate students, Ph.D. thesis, University of Pittsburgh (2015).
- Zhu and Singh [2012] G. Zhu and C. Singh, Surveying students’ understanding of quantum mechanics in one spatial dimension, American Journal of Physics 80, 252 (2012).
- Cataloglu and Robinett [2002] E. Cataloglu and R. Robinett, Testing the development of student conceptual and visualization understanding in quantum mechanics through the undergraduate career, American Journal of Physics 70, 238 (2002).
- Wuttiprom et al. [2009] S. Wuttiprom, M. D. Sharma, I. D. Johnston, R. Chitaree, and C. Soankwan, Development and use of a conceptual survey in introductory quantum physics, International Journal of Science Education 31, 631 (2009).
- Allain [2001] R. J. Allain, Investigating the relationship between student difficulties with the concept of electric potential and the concept of rate of change (North Carolina State University, 2001).
- Aslanides and Savage [2013] J. Aslanides and C. M. Savage, Relativity concept inventory: Development, analysis, and results, Physical Review Special Topics – Physics Education Research 9, 010118 (2013).
- Nieminen et al. [2010] P. Nieminen, A. Savinainen, and J. Viiri, Force concept inventory-based multiple-choice test for investigating students’ representational consistency, Physical Review Special Topics – Physics Education Research 6, 020109 (2010).
- Mashood and Singh [2012] K. Mashood and V. A. Singh, An inventory on rotational kinematics of a particle: unravelling misconceptions and pitfalls in reasoning, European Journal of Physics 33, 1301 (2012).
- Suárez et al. [2023] M. Suárez, S. Pandiella, and J. Benegas, Tutorials+ PhET: a simple and efficient active-learning approach for the teaching of kinematics of circular motion in a technically-oriented high school, Physics Education 58, 035005 (2023).
- Rimoldini and Singh [2005] L. G. Rimoldini and C. Singh, Student understanding of rotational and rolling motion concepts, Physical Review Special Topics – Physics Education Research 1, 010102 (2005).
- Singh [2006] C. Singh, Student understanding of symmetry and Gauss’s law of electricity, American journal of physics 74, 923 (2006).
- Bailey et al. [2012] J. M. Bailey, B. Johnson, E. E. Prather, and T. F. Slater, Development and validation of the star properties concept inventory, International Journal of Science Education 34, 2257 (2012).
- Brown and Singh [2021] B. Brown and C. Singh, Development and validation of a conceptual survey instrument to evaluate students’ understanding of thermodynamics, Physical Review Physics Education Research 17, 010104 (2021).
- Yeo and Zadnik [2001] S. Yeo and M. Zadnik, Introductory thermal concept evaluation: Assessing students’ understanding, The Physics Teacher 39, 496 (2001).
- Wattanakasiwich et al. [2013] P. Wattanakasiwich, P. Taleab, M. D. Sharma, and I. D. Johnston, Construction and implementation of a conceptual survey in thermodynamics, International Journal of Innovation in Science and Mathematics Education 21 (2013).
- Slater [2014] S. J. Slater, The development and validation of the test of astronomy standards (TOAST)., Journal of Astronomy & Earth Sciences Education 1, 1 (2014).
- Klein et al. [2020] P. Klein, A. Lichtenberger, S. Küchemann, S. Becker, M. Kekule, J. Viiri, C. Baadte, A. Vaterlaus, and J. Kuhn, Visual attention while solving the test of understanding graphs in kinematics: an eye-tracking analysis, European Journal of Physics 41, 025701 (2020).
- Barniol and Zavala [2014] P. Barniol and G. Zavala, Test of understanding of vectors: A reliable multiple-choice vector concept test, Physical Review Special Topics-Physics Education Research 10, 010121 (2014).
- Microsoft [2024] Microsoft, Azure AI Services, https://azure.microsoft.com/en-us/products/ai-services (accessed June 2024).
- Danilák [2024] M. Danilák, langdetect, https://pypi.org/project/langdetect/ (accessed December 2024).
- OpenAI [2025] OpenAI, Introducing GPT-o1, https://openai.com/o1/ (accessed January 2025).